40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 21 00:45:30 2015, Koji, Update, IOO, IMC fix PDRFsignal_cable.JPGQsignal_cable1.JPGQsignal_cable2.JPGQsignal_cable3.JPG
    Reply  Sat Nov 21 00:46:33 2015, Koji, Update, IOO, IMC OLTF IMC_OLTF.pdfPMC_noise_comparison.pdf
       Reply  Sun Nov 22 12:12:17 2015, Koji, Update, IOO, IMC OLTF 
    Reply  Sun Nov 22 07:09:01 2015, rana, Update, IOO, IMC fix 
       Reply  Sun Nov 22 12:07:09 2015, Koji, Update, IOO, IMC fix 
          Reply  Mon Nov 23 22:12:10 2015, Koji, Update, IOO, LO level check for the IMC demod board IMG_2137.JPGIMG_2138.JPG
             Reply  Wed Nov 25 04:24:21 2015, rana, Update, IOO, LO level check for the IMC demod board 
                Reply  Wed Nov 25 10:07:10 2015, Koji, Update, IOO, LO level check for the IMC demod board 
                   Reply  Wed Nov 25 16:46:10 2015, Koji, Update, IOO, LO level check for the IMC demod board D990511-40m_151123.pdf
                      Reply  Wed Nov 25 20:07:35 2015, rana, Update, IOO, LO level check for the IMC demod board 
                   Reply  Mon Nov 30 14:12:14 2015, ericq, Update, LSC, LO level check for the LSC RF distribution box 
                      Reply  Mon Nov 30 15:17:57 2015, Koji, Update, LSC, LO level check for the LSC RF distribution box 
Message ID: 11795     Entry time: Sat Nov 21 00:46:33 2015     In reply to: 11794     Reply to this: 11798
Author: Koji 
Type: Update 
Category: IOO 
Subject: IMC OLTF 

Here is the comparison before and after the fix.

Before the work, the UGF was ~40kHz. The phase margin was ~5deg. This caused huge bump of the frequency noise.

After the LO power increase, I had to reduce the MC loop gain (VCO Gain) from 18dB to 6dB. This resulted 4dB (x2.5) increase of the OLTF. This means that my fix increased the optical gain by 16dB (x6.3). The resulting UGF and phase mergin were measured to be 117kHz and 31deg, respectively.

 

Now I was curious to see if the PMC err shows reasonable improvement when the IMC is locked. Attachment 2 shows the latest comparison of the PMC err with and without the IMC locked. The PMC error has been taken up to 500kHz. The errors were divided by 17.5kHz LPF and 150kHz LPF to compensate the sensing response. The PMC cavity pole was ignored in this calculation. T990025 saids the PMC finesse is 4400 and the cavity pole is 174kHz. If this is true, this also needs to be applied.

Observations:

1. Now we can see improvement of the PMC error in the region between 10kHz to 70kHz.

2. The sharp peak at 8kHz is due to the marginally stable PMC servo. We should implement another notch there. T990025 suggests that the body resonance of the PMC spacer is somewhere around there. We might be able to damp it by placing a lossy material on it.

3. Similarly, the features at 12kHz and 28kHz is coming from the PMC. They are seen in the OLTF of the PMC loop.

4. The large peak at 36kHz does not change with the IMC state. This does mean that it is coming from the laser itself, or anything high-Q of the PMC. This signal is seen in the IMC error too.

5. 72kHz, 108kHz, 144kHz: Harmonics of 36kHz?

6. Broad feature from 40kHz to 200kHz. The IMC loop is adding the noise. This is the frequency range of the PC drive. Is something in the PC drive noisy???
 

7. The feature at 130kHz. Unknown. Seems not related to IMC. The laser noise or the PMC noise.


Remaining IMC issues:

Done (Nov 23, 2015) - 29.5MHz oscillator output degraded. Possibly unstable and noisy. Do we have any replacement? Can we take a Marconi back from one of the labs?

Done (Nov 23, 2015) - Too high LO?

- Large 36kHz peak in the IMC

- IMC loop shape optimization

- IMC locking issue. The lock streatch is not long.

- IMC PC drive issue. Could be related to the above issue.

Maybe not relevant - PC drive noise?

Attachment 1: IMC_OLTF.pdf  71 kB  Uploaded Sat Nov 21 15:27:56 2015  | Hide | Hide all
IMC_OLTF.pdf
Attachment 2: PMC_noise_comparison.pdf  55 kB  Uploaded Sat Nov 21 15:28:09 2015  | Hide | Hide all
PMC_noise_comparison.pdf
ELOG V3.1.3-